All of these problems involve preparation and use

I just need to figure out number 3 and 4. thank you!!
Answer #1

Data.

t1/2 = 0.693/k

k = rate/[A]

--

[A] = 20 ng/L

t1/2 = 8.4 hrs

--

Answer.

3. From half life equation, we know that a first order reaction is independent from initial concentration, this is a measure that approximates the magnitude of a rate constant, the lower the half life, and the greater will be K.

t1/2 = 0.693/k

@ 25°C = 298 K

30240 s = 0.693/k

k1 = 2.29x10-5 s-1

--

@ 4°C = 277 K

41400 s = 0.693/k

k2 = 1.67x10-5 s-1

--

The dependent of a rate constant regard to temperature is expressed by Arrhenius's equation:

k = Ae-Ea/RT

Where;

Ea = activation energy (KJ/mol)

A = frequency factor

From this equation we can calculate the relation between two rate constant at different temperatures. Applying ln in both sides:

lnk1 = lnA - Ea/RT1

lnk2 = lnA - Ea/RT2

by subtracting lnk2 from lnk1:

ln k1/k2 = Ea/R(T1-T2/T1T2)

ln(2.29x10-5 s-1/1.67x10-5 s-1) = Ea/8.314 J/K.mol (298 - 277/298*277)

Ea = 44845.7958 J/mol = 44.85 KJ/mol

--

4. If vaccine is prepared at 8:30 and will be used until 12:00 then we have 3 hours and a half (11880 s) for its usage. From rate equation we can obtain:

rate = k[A]

-Δ[A]/Δt = k[A]

From here we can demonstrate:

ln[A]/[Ao] = -kt

ln[A] = -kt + ln [Ao]

ln[A] = -(1.67x10-5 s-1)(11880 s) + ln(20 ng/L)

eln[A] = e2.7986

[A] = 16.4228 ng/mol

--

The vaccine should not be fully effective because the concentration is lower than the ideal dosage.